“Blast it! said David Wilson, president of Teledex Company. “Weve just lost the bid on the Koopers job by $3,000. It seemsb. Determine the amount of manufacturing overhead cost that would have been applied to the Koopers job. Manufacturing overhea
Answer #1

1a) Predetermine overhead rate = 844200/603000 = 140% of labor cost

1b) Manufacturing overhead applied = 9900*1.4 = 13860

2a) Predetermine overhead rate :

Fabricating department = 351750/201000 = 175% of labor cost

Machining department = 402000/100500 = 400% of labor cost

Assembly department = 90450/301500 = 30% of labor cost

2b) Applied overhead = (3000*1.75+600*4+6300*30%) = 9540